# Entanglement Microscopy: Tomography and Entanglement Measures via Quantum Monte Carlo

### v3

@@ Abstract
We develop a protocol, dubbed entanglement microscopy, to obtain the full reduced density matrix associated with subregions in quantum Monte Carlo simulations for bosonic and fermionic many-body systems. Our microscopy allows to perform quantum state tomography, and thus gives access to true entanglement measures, such as the logarithmic negativity (LN). We exemplify our method by studying the phase diagram near quantum critical points (QCP) in 2 spatial dimensions: the transverse field Ising model and a Gross-Neveu-Yukawa transition of Dirac fermions. Our main results are: i) the Ising QCP exhibits short-range entanglement with a finite sudden death of the LN both in space and temperature; ii) the Gross-Neveu QCP has a power-law decaying fermionic LN consistent with conformal field theory (CFT) exponents; iii) going beyond bipartite entanglement, we find no detectable 3-party entanglement by our two witnesses in a large parameter window near the Ising QCP in 2d, in contrast to 1d. We also analytically obtain the large-temperature power-law scaling of the fermionic LN for general interacting systems. Our approach allows one to perform quantum state tomography locally in a way that is analogous to atomic-scale imaging with a scanning tunneling microscope. Controlled entanglement microscopy opens a new window into quantum matter, with countless systems waiting to be explored.
@@
Quantum entanglement can reveal the fundamental organizing principles of quantum matter, ranging from non-Fermi liquids in high-temperature superconductors to black holes [1, 2]. However, obtaining the entanglement structure of realistic systems is a challenging task. When dealing with physically relevant situations that include finite temperature or the study of microscopic subregions, one has to work with mixed states (density matrices), and simple quantities like the von Neumann or Rényi entropies cannot be used as measures of entanglement. In the simplest case of bipartite entanglement, one can for example exploit the partial transpose to construct the (logarithmic) negativity [3, 4, 5, 6], which is a true entanglement measure not polluted by classical correlations. However, it is challenging to compute in many-body systems for subregions with many degrees of freedom (including quantum field theories) where one can only obtain related quantities via the replica trick [7, 8]. These Rényi analogs are not entanglement measures and thus cannot be used to obtain the entanglement structure. A striking example is in boson/spin systems poised near a quantum critical point (QCP), where the logarithmic negativity (LN) decays faster than any power, as shown for CFTs in 1d [8] and higher $d$ [9], while the Rényi moments decay algrebraically. These are special cases of a general phenomenon regarding the “fate of entanglement” [10] where any multi-partite form of entanglement (including but more general than bipartite) entirely disappear as a system heats up, evolves in time, or as its parts grow separated.
To reveal the true microscopic structure of entanglement, in a way analogous to what the scanning tunneling microscope does for imaging atomic-scale densities, we use quantum Monte Carlo (QMC) simulations to perform full tomography on the state of microscopic subregions. More precisely, we use a path-integral QMC approach that allows us to directly sample the reduced density matrix (RDM) of various subregions, as shown in Fig. 1. Whereas a scanning tunneling microscope usually works with one tip, here we exploit numerous “tips” to capture $m$-party entanglement. In the present paper, we illustrate such entanglement microscopy for $m=2,3$.
@@
Specifically, we use entanglement microscopy to scrutinize systems near generic 2d quantum phase transitions: first the celebrated QCP in the transverse field Ising model (TFIM) on the square lattice, and second a Gross-Neveu-Yukawa QCP of interacting Dirac fermions. Regarding bipartite entanglement, our tomography allows the computation of the LN in the entire phase diagram including at finite temperature with the main findings: i) the Ising model exhibits short-range entanglement with finite sudden death as a function of space and temperature; ii) the fermionic variant of the LN, which takes into account the parity superselection rule, shows power-law decay at the Dirac QCP for large separations with exponents consistent with the conformal field theory (CFT) prediction [9]. iii) Going beyond bipartite entanglement, we do not detect any genuine 3-party entanglement in a large window near the 2d Ising QCP, but find its presence in the 1d model both in QMC and with the exact solution via fermionization. Overall, the 2d Ising model shows less bipartite and tripartite entanglement compared to 1d [11, 12, 13, 14];
Our entanglement microscopy scheme offers advantages in terms of immense versatility and computational efficiency. By focusing on microscopic or so-called “skeletal” subregions [15], entanglement microscopy not only gives the full RDM and true entanglement measures, but it does so without the need of replicating the system [16], which leads to increases in computational complexity, numerical instabilities even with advanced QMC algorithms for entanglement entropy calculation [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], and stringent requirement of data quality during the analytic continuation procedures in the entanglement spectrum computation [28, 29].
@@
Figure 1: Entanglement microscopy: sampling the reduced density matrix in QMC. Panel (a) shows the partition function in the space ($r$) - time ($\tau$) manifold appearing in usual QMC simulations, with the periodic structure of imaginary time. Panel (b) shows how the RDM is sampled in such QMC simulations, where one imposes open boundary conditions in time for the skeletal sites in subregion $A$. From the QMC histogram of configurations of these sites, in this case, there will be $4\times 4$ configurations for the 2 spins in $A$, the corresponding matrix elements in the RDM are recorded. The lower panel in (b) demonstrates the spatial partition of $A$ and $B$ in such a setting.
@@ Results
Entanglement microscopy.
The entanglement microscopy protocol is illustrated in Fig. 1, in the realization of the path-integral QMC for quantum spin systems. In such stochastic series expansion (SSE) QMC simulations [30], the RDM of the spins in subregion $A$, $\rho_{A}$, is evaluated as
@@
$\begin{split}\rho_{A}&=\frac{1}{Z}\operatorname{Tr}_{B}\{e^{-\beta H}\}=\frac{1}{Z}\sum_{\xi}\bra{\xi}e^{-\beta H}\ket{\xi}\\ &=\frac{1}{Z}\sum_{\{b_{1},b_{2}\cdots b_{M}\}}\frac{\beta^{n}(M-n)!}{M!}\sum_{\xi}\bra{\xi}H_{b_{M}}\cdots H_{b_{2}}H_{b_{1}}\ket{\xi},\end{split}$ (1)
where {|$\{\ket{\xi}\}$ forms a complete basis of the environment $B$. The sampling of the RDM is done in an enlarged ensemble where one does not enforce the left state and right state to be the same in region $A$, i.e., we discard the periodic boundary condition in time. As shown in Fig. 1 (b), by keeping the periodic boundary condition in region $B$ and open boundary in region $A$, in each Monte Carlo step, one records the spin configuration on each site in region $A$, and the frequency of each element appearing is proportional to its corresponding element in the RDM.
For interacting fermions, one employs the determinant quantum Monte Carlo (DQMC) scheme [31, 32, 33]. In DQMC, the sampling of the RDM is attributed to the fact that fermions are considered free in each auxiliary field configuration, thus the interacting RDM is the weighted average of free RDMs [18]. For the case where region $A$ only contains two sites, the interacting RDM in the occupation-number basis can be constructed with the density correlation function, and fermion Green’s function [9].
We note that DQMC calculation of Rényi negativity with bipartition at finite temperature of interacting fermions, via the replica trick, has recently been worked out [34]. However, since the RDM therein is not explicitly computed, only integer moments of LN can be obtained.
@@
Figure 2: Logarithmic negativity for adjacent sites. Color code plots of (a) $\mathcal{E}^{b}$ in 2d TFIM simulated in a $40\times 40$ square lattice with $\beta$ up to 10, and (b) $\mathcal{E}^{f}$ in 2d fermion $t-V$ model simulated in a $6\times 6$ square lattice with $\beta$ up to 15. In both panels, the red curves represent the finite-temperature phase boundaries which end at a zero temperature QCP located at $h_{c}=3.044$ in 2d TFIM [35] and $V_{c}=1.27$ in $t-V$ model. The purple dashed lines show the maxima of $\mathcal{E}^{b/f}$ for each temperature. In both models, the maxima have converged with system size, as is shown in the SM [36]. The bottom panel illustrates the schematic phase diagrams.
Entanglement measures & Multipartite entanglement criteria. A strong advantage of our entanglement microscopy is the versatility of the full RDM: we can use it to compute not only the expectation value of any observable but also any entanglement measure involving the degrees of freedom in $A$. An example is the logarithmic negativity that can be used to shed light on bipartite entanglement, and its evolution in the entire phase diagram [15, 9]. More interestingly, we can also evaluate various multipartite entanglement criteria, and these reveal more collective forms of entanglement.
Bipartite entanglement. The logarithmic negativity is calculated with the RDM $\rho_{A}$ where region $A$ is subsequently divided into two part, $A_{1}$ and $A_{2}$, and one takes the partial transpose in region $A_{1}$. It is defined as [3, 4, 5]
@@
Tripartite entanglement. In the above, we have discussed bipartite entanglement between two subregions. Entanglement microscopy allows us to delve into the much richer realm of multi-party entanglement. We illustrate this by studying tripartite entanglement between 3 sites in the 2d quantum Ising model. In particular, we shall search for genuine tripartite entanglement (GTE), which cannot be described in terms of 2-party entanglement between any two subgroups of spins. A state with GTE cannot be written as a mixture of bi-separable physical density matrices, $\rho\neq\sum_{k}p_{k}\rho_{A}^{k}\rho_{BC}^{k}+q_{k}\rho_{AB}^{k}\rho_{C}^{k}+r_{k}\rho_{AC}^{k}\rho_{B}^{k}$, where $p_{k},q_{k},r_{k}$ are non-negative. A simple example of GTE is given by a pure cat state (also known as GHZ state) of 3 spins, $a|000\rangle+b|111\rangle$. Criteria to detect GTE are more involved than those of the simple bipartite measures such as the LN. However, some criteria exist that can detect GTE by establishing that a state is not bi-separable. Here, we shall focus on two such separability criteria, that we call $W_{1}$ and $W_{2}$ [41]. Let us describe $W_{1}$ ($W_{2}$ is analogously defined, see Supplemental Material (SM) [36]):
@@
where $\rho_{i,j}$ are the matrix element of the 3-site RDM. In order to get a meaningful answer, one has to remove the basis dependence by maximizing among all local unitary transformations acting on the RDM. Being local, such transformations leave the GTE structure invariant. If $W_{i}>0$, there is GTE, while no conclusion can be reached if $W_{i}\leq 0$. Although not sensitive to all forms of GTE, the $W_{i}$ criteria do tend to perform well in a variety of states, especially when used together [41].
Figure 3: Spatial and temporal dependence of the logarithmic negativity. (a) The decay of the two-site $\mathcal{E}^{b}$ and $\mathcal{E}^{b}_{3}$ as a function of separation $r$ at the QCP $h_{c}=3.04$. The simulation is done with $L=\beta=40$. $\mathcal{E}^{b}$ exhibits a sudden death beyond one lattice constant, and the $\mathcal{E}^{b}_{3}$ exhibits a power-law decay with the fit shown as a purple dashed line; the exponent $3.9(2)$ is consistent with the expected Ising CFT value $4\Delta_{\sigma}+2\approx 4.073$ [9, 42]. (b) The decay of $\mathcal{E}^{f}$ versus separation $r$ in the $t-V$ model with various system sizes, at the QCP $V_{c}=1.27$. The simulation is done with $L=\beta$ going from 6 to 14. The blue dashed line shows the algebraic scaling with the power predicted by CFT analysis [9, 43], with the expected GNY scaling dimension $4\Delta_{\psi}\approx 4.174$. Panel (c) and (d) demonstrate the temperature dependence of $\mathcal{E}^{b}$ for the $L=40$ TFIM, and $\mathcal{E}^{f}$ for the $L=6$ $t-V$ model at various values of $h$ and $V$. In (c) the $\mathcal{E}^{b}$ always exhibit a finite sudden death temperature and in (d) $\mathcal{E}^{f}$ decays as $1/T^{2}$ for adjacent sites $r=1$, while $1/T^{6}$ for sites separated by $r=3$ bonds.
@@
it consists of nearest neighbor hopping term with $\pi$-flux and nearest neighbor repulsion. We set $t=1$ as the energy unit. The $\pi$-flux hopping term on the square lattice ($\theta_{i,j}=\frac{\pi}{2}$ on every other column of vertical bonds and 0 otherwise) leads to a dispersion with two Dirac cones at the $X=(\pi,\pm\frac{\pi}{2})$ points in the Brillouin zone (BZ) taking two sites horizontally as a unit cell. At half filling, a positive $V$ drives the system from a Dirac semi-metal (DSM) to a $\mathbb{Z}_{2}$-symmetry spontaneously broken charge density wave (CDW) state, where one sublattice acquires a greater density compared to the other. The QCP is described by the “chiral Ising” Gross-Nevey-Yukawa (GNY) theory with global symmetry $O(N_{f})^{2}\rtimes\mathbb{Z}_{2}$ and in our case $N_{f}=2$ two-component Dirac fermions. The position of the QCP, the critical exponents of the Gross-Neveu transition, have been systematically worked out using QMC, field-theoretical expansions, and conformal bootstrap, for both $N_{f}=2$ [44, 43, 45, 46, 47, 48], and other values of $N_{f}$ [49, 43, 45, 46, 50, 47, 48, 51, 52, 53, 42, 54, 55, 56]. The schematic $T-V$ phase diagram is shown in lower part of the panel (b) in Fig. 2.
@@
By simulating this model with $L=\beta=6,8,10,12,14$, we can probe entanglement near the chiral GNY-Ising QCP. We located the QCP at $V_{c}=1.27$ via finite-size scaling and using the critical exponents from conformal bootstrap [43]. The detailed QMC finite-size data analysis and collapse are shown in the SM [36].
Logarithmic negativity. We now present the results for $\mathcal{E}^{b}$ of the TFIM and $\mathcal{E}^{f}$ of the $t-V$ model.
@@
Ising model. Fig. 2 (a) shows the nearest neighbor $\mathcal{E}^{b}$ at finite temperature for the square lattice TFIM of linear system size $L=40$ with inverse temperature $\beta$ up to 10. The maxima of $\mathcal{E}^{b}$ at each temperature, indicated by the purple dashed line, go from the disordered state at high temperature to $h=3.1$ at the lowest temperature $T=1/10$, which is close to but distinct from the QCP, $h_{c}=3.044$. At all $h$ values, $\mathcal{E}^{b}$ has a sudden-death temperature, beyond which $\mathcal{E}^{b}$ is identically zero corresponding to the dark blue region in the phase diagram; it is also shown in panel (c) of Fig. 3. The existence of a thermal sudden death follows from the general fact that the infinite temperature state lies well within the continent of separable states, and so entanglement must disappear at a finite temperature corresponding to the point at which the system enters the continent [10].
@@
Let us now examine the fate of entanglement as the 2 sites become progressively more separated. Fig. 3 (a) shows the $\mathcal{E}^{b}$ between two sites separated with bond-distance $r=|i-j|$. $\mathcal{E}^{b}$ drops to 0, within machine precision, even for the next-nearest neighbor separation, $r=\sqrt{2}$. Strikingly, the 2d TFIM possesses the shortest possible range of entanglement for two spins, as it only exists for nearest neighbor, even shorter than the 1d case where the sudden-death distance was found to be $r_{\textrm{sd}}=3$ [57, 12, 14, 36]. Such a sudden death separation again follows since the state at large separations lies within the separable continent [10].
@@
We also computed the third order moment $\mathcal{E}^{b}_{3}$ at the Ising QCP. The moments of the LN are defined [7, 8] as $\mathcal{E}^{b}_{n}=\ln\frac{\operatorname{Tr}\rho_{A}^{n}}{\operatorname{Tr}[(\rho_{A}^{T_{1}})^{n}]},$ which can be easily computed with the RDM obtained via microscopy, or one can sample it by using standard QMC and exploiting the replica trick [58]. The orange line in the inset of Fig. 3 (a) shows the power-law decay of $\mathcal{E}^{b}_{3}$ with fitted power 3.9(2), which agrees with the CFT prediction $4\Delta_{\sigma}+2\approx 4.073$ [9, 42]. The power-law decay of $\mathcal{E}^{b}_{3}$ is in sharp contrast with the rapid sudden death of $\mathcal{E}^{b}$, and arises since the former is not a proper entanglement measure, and is thus polluted by non-entangling correlations.
@@
Fermion $t-V$ model. Fig. 2 (b) shows the analogous results for $\mathcal{E}^{f}$ for the square lattice $t-V$ model of linear system size $L=6$ with $\beta$ up to 15. As denoted by the purple dashed line, the maxima starts in the ordered CDW phase at high temperature and converges to $V\approx 1$ at low temperature, which is in the DSM phase. It roughly follows the finite-temperature phase boundary, obtained from finite-size scaling [36] and indicated by the red line, but clearly differs from it at most temperatures. At high $T$, the nearest neighbor $\mathcal{E}^{f}$ does not suffer a sudden death. Instead, it decays algebraically as shown in Fig. 2 (b) and Fig. 3 (d).

### v4

# Entanglement Microscopy: Tomography and Entanglement in Many-Body Systems

@@ Abstract
We employ a protocol, dubbed entanglement microscopy, to reveal the multipartite entanglement encoded in the full reduced density matrix of microscopic subregion both in spin and fermionic many-body systems. We exemplify our method by studying the phase diagram near quantum critical points (QCP) in 2 spatial dimensions: the transverse field Ising model and a Gross-Neveu-Yukawa transition of Dirac fermions. Our main results are: i) the Ising QCP exhibits short-range entanglement with a finite sudden death of the LN both in space and temperature; ii) the Gross-Neveu QCP has a power-law decaying fermionic LN consistent with conformal field theory (CFT) exponents; iii) going beyond bipartite entanglement, we find no detectable 3-party entanglement with our two witnesses in a large parameter window near the Ising QCP in 2d, in contrast to 1d. We further establish the singular scaling of general multipartite entanglement measures at criticality, and present an explicit analysis in the tripartite case. We also analytically obtain the large-temperature power-law scaling of the fermionic LN for general interacting systems. Entanglement microscopy opens a rich window into quantum matter, with countless systems waiting to be explored.
Quantum entanglement can reveal the fundamental organizing principles of quantum matter, ranging from non-Fermi liquids in high-temperature superconductors to black holes [1, 2]. However, extracting the entanglement structure in realistic many-body systems remains a significant challenge. A key step toward understanding these systems is reconstructing the quantum state of microscopic subregions, namely those that contain a small number of sites. By focusing on smaller subregions, we can fully reconstruct the reduced density matrix (RDM), potentially giving us access to complete entanglement information. The entanglement microscopy protocol thus involves first obtaining the RDM numerically, and then calculating multipartite entanglement within these subregions, providing deeper insights into quantum critical systems.
@@
This approach complements existing methods that focus on extended ensembles [3, 4, 5, 6, 2, 7, 8, 9, 10, 11, 12], which handle larger subregions but are often limited in the choice of entanglement measures. It also differs from global measures like Quantum Fisher Information (QFI) and Quantum Variance (QV) which capture coarser quantum correlations in the entire system [13, 14]. In the simplest case of bipartite entanglement, having the full RDM allows us to compute the (logarithmic) negativity [15, 16, 17, 18], an entanglement monotone not polluted by classical correlations. However, negativity is challenging to compute in many-body systems for subregions with many degrees of freedom (including quantum field theories) where one can only obtain related quantities via the replica trick [19, 20]. These Rényi analogs are not entanglement monotones and thus cannot be used to obtain the entanglement structure. This limitation becomes especially clear in boson/spin systems poised near a quantum critical point (QCP), where the logarithmic negativity (LN) decays faster than any power, as shown for CFTs in 1d [20] and higher $d$ [21], while the Rényi moments decay algebraically. These are special cases of a general phenomenon regarding the “fate of entanglement” [22] where any multi-partite form of entanglement (including but more general than bipartite) entirely disappears as a system heats up, evolves in time during certain protocols, or as its parts grow separated.
@@
In the first step of our protocol, we compute the reduced density matrix (RDM) for microscopic or “skeletal” subregions of many-body systems [23]. This can be achieved through various computational approaches such as Exact Diagonalization (ED), Quantum Monte Carlo (QMC), or tensor network approaches such as the Density Matrix Renormalization Group (DMRG). In our work, we focus on numerically constructing the RDM using ED and QMC [24]. To be more specific, we scrutinize systems near generic 2d quantum phase transitions, including (1) the celebrated QCP in the transverse field Ising model (TFIM) on the square lattice and (2) a Gross-Neveu-Yukawa (GNY) QCP of interacting Dirac fermions. Regarding bipartite entanglement, our microscopy yields the LN in the entire phase diagram including at finite temperature with the main findings: i) the Ising model exhibits short-range entanglement with finite sudden death as a function of space and temperature; ii) the fermionic variant of the LN, which takes into account the parity superselection rule, shows power-law decay at the GNY QCP for large separations with exponents consistent with the conformal field theory (CFT) prediction [21]. iii) Going beyond bipartite entanglement, we do not detect any genuine 3-party entanglement in a large window near the 2d Ising QCP, but find its presence in the 1d model both in QMC and with the exact solution via fermionization. Overall, the 2d Ising model shows less bipartite and tripartite entanglement compared to 1d [25, 26, 27, 28].
@@
Figure 1: Sampling the reduced density matrix in path-integral QMC. Panel (a) shows the partition function in the space ($r$) - time ($\tau$) manifold appearing in usual QMC simulations, with the periodic structure of imaginary time. Panel (b) shows how the RDM is sampled in such QMC simulations, where one imposes open boundary conditions in time for the skeletal sites in subregion $A$. From the QMC histogram of configurations of these sites, in this case, there will be $4\times 4$ configurations for the 2 spins in $A$, the corresponding matrix elements in the RDM are recorded. The lower panel in (b) demonstrates the spatial partition of $A$ and $B$ in such a setting.
@@ Results
Entanglement microscopy
To extract the entanglement information in microscopic subregions for spin models, we use the algorithm introduced in Ref. [24], schematically shown in Fig. 1. In general, the matrix elements of the RDM for spins in subregion $A$, $\langle\alpha|\rho_{A}|\tilde{\alpha}\rangle$ is expressed as
@@
$\begin{split}\langle\alpha|\rho_{A}|\tilde{\alpha}\rangle&=\frac{1}{Z}\bra{\alpha}\operatorname{Tr}_{B}e^{-\beta H}\ket{\tilde{\alpha}}\\ &=\frac{1}{Z}\sum_{\xi}\bra{\alpha,\xi}e^{-\beta H}\ket{\tilde{\alpha},\xi},\end{split}$ (1)
where $\{|\xi\rangle\}$ forms a complete basis for the environment region $B$ and $\beta$ is the inverse temperature. Eq. (1) can be viewed as a modified partition function with imaginary-time periodicity only in the region $B$ [24], which can be easily simulated in path-integral QMC as shown in Fig. 1. Please refer to the latter section for a detailed description.
For interacting fermions, one employs the determinant quantum Monte Carlo (DQMC) scheme [29, 30, 31]. In DQMC, the sampling of the RDM is attributed to the fact that fermions are considered free in each auxiliary field configuration, thus the interacting RDM is the weighted average of free RDMs [32]. Recent advances in DQMC have enabled the reconstruction of the RDM for small subregions in many-body quantum systems, such as a four-site plaquette in the two-dimensional fermion Hubbard model [33]. However, in that work, genuine entanglement between sites within the plaquette has not been studied. Similarly, studies using dynamical mean-field theory (DMFT) on the Kagome Hubbard model have focused on subregions and examined only some diagonal elements of the RDM [34], without exploring genuine entanglement. For the case that we focus on, where region $A$ contains two sites, the interacting RDM in the occupation-number basis can be constructed with the density correlation function, and fermion Green’s function [35]. We also note that DQMC calculation of Rényi negativity with bipartition at finite temperature of interacting fermions, via the replica trick, has recently been worked out [36]. However, since the RDM therein is not explicitly computed, only integer moments of LN can be obtained.
@@
Figure 2: Logarithmic negativity for adjacent sites. Color code plots of (a) $\mathcal{E}^{b}$ in 2d TFIM simulated in a $40\times 40$ square lattice with $\beta$ up to 40, and (b) $\mathcal{E}^{f}$ in 2d fermion $t-V$ model simulated in a $6\times 6$ square lattice with $\beta$ up to 15. In both panels, the red curves represent the finite-temperature phase boundaries which end at a zero temperature QCP located at $h_{c}=3.044$ in 2d TFIM [37] and $V_{c}=1.27$ in $t-V$ model. The purple dashed lines show the maxima of $\mathcal{E}^{b/f}$ for each temperature. In both models, the maxima have converged with system size, as is shown in the SM [38]. The bottom panel illustrates the schematic phase diagrams.
Bipartite entanglement. Bipartite entanglement quantifies the quantum correlations between two subsystems, which cannot be reduced to classical correlations. Widely used bipartite entanglement measures include entanglement entropy, concurrence, and logarithmic negativity, the latter valued for its simplicity and applicability to both pure and mixed states, making it a robust tool for studying entanglement in systems near quantum critical points [39]. Consider a region $A$, which is divided into two parts, $A_{1}$ and $A_{2}$. The logarithmic negativity is defined as a function of the partial transpose on the RDM $\rho_{A}$ in region $A_{1}$ [15, 16, 17]
@@
Tripartite entanglement. In the above, we have discussed bipartite entanglement between two subregions. Entanglement microscopy can also delve into the much richer realm of multi-party entanglement. We illustrate this by studying tripartite entanglement between 3 sites in the 2d quantum Ising model. In particular, we shall search for genuine tripartite entanglement (GTE), which cannot be described in terms of 2-party entanglement between any two subgroups of spins. A state with GTE cannot be written as a mixture of bi-separable physical density matrices, $\rho\neq\sum_{k}p_{k}\rho_{A}^{k}\rho_{BC}^{k}+q_{k}\rho_{AB}^{k}\rho_{C}^{k}+r_{k}\rho_{AC}^{k}\rho_{B}^{k}$, where $p_{k},q_{k},r_{k}$ are non-negative. A simple example of GTE is given by a pure cat state (also known as GHZ state) of 3 spins, $a|000\rangle+b|111\rangle$, with non-zero $a,b$. Criteria to detect GTE are more involved than those of the simple bipartite measures such as the LN. However, some criteria exist that can detect GTE by establishing that a state is not bi-separable. Here, we shall focus on two such separability criteria, that we call $W_{1}$ and $W_{2}$ [44]. Let us describe $W_{1}$ ($W_{2}$ is analogously defined, see Supplemental Material (SM) [38]):
@@
where $\rho_{i,j}$ are the matrix element of the 3-site RDM. In order to get a meaningful answer, one has to remove the basis dependence by maximizing among all local unitary (LU) transformations acting on the RDM, $\rho\to U\rho U^{\dagger}$ with $U=U_{1}\otimes U_{2}\otimes U_{3}$. Being local, such transformations leave the GTE structure invariant. If $W_{i}>0$, there is GTE, while no conclusion can be reached if $W_{i}\leq 0$. Although not sensitive to all forms of GTE, the $W_{i}$ criteria do tend to perform well in a variety of states, especially when used together [44].
Figure 3: Spatial and temporal dependence of the logarithmic negativity. (a) The decay of the two-site $\mathcal{E}^{b}$ and $\mathcal{E}^{b}_{3}$ as a function of separation $r$ at the QCP $h_{c}=3.04$. The simulation is done with $L=\beta=40$. $\mathcal{E}^{b}$ exhibits a sudden death beyond one lattice constant, and the $\mathcal{E}^{b}_{3}$ exhibits a power-law decay with the fit shown as a purple dashed line; the exponent $3.9(2)$ is consistent with the expected Ising CFT value $4\Delta_{\sigma}+2\approx 4.073$ [21, 45]. (b) The decay of $\mathcal{E}^{f}$ versus separation $r$ in the $t-V$ model with various system sizes, at the QCP $V_{c}=1.27$. The simulation is done with $L=\beta$ going from 8 to 16. The blue dashed line shows the algebraic scaling with the power predicted by CFT analysis [21, 46], with the expected GNY scaling dimension $4\Delta_{\psi}\approx 4.174$. Panel (c) and (d) demonstrate the temperature dependence of $\mathcal{E}^{b}$ for the $L=40$ TFIM, and $\mathcal{E}^{f}$ for the $L=6$ $t-V$ model at various values of $h$ and $V$. In (c) the $\mathcal{E}^{b}$ always exhibit a finite sudden death temperature and in (d) $\mathcal{E}^{f}$ decays as $1/T^{2}$ for adjacent sites $r=1$, while $1/T^{6}$ for sites separated by $r=3$ bonds.
@@
it consists of nearest neighbor hopping term with $\pi$-flux and nearest neighbor repulsion. We set $t=1$ as the energy unit. The $\pi$-flux hopping term on the square lattice ($\theta_{i,j}=\frac{\pi}{2}$ on every other column of vertical bonds and 0 otherwise) leads to a dispersion with two Dirac cones at the $X=(\pi,\pm\frac{\pi}{2})$ points in the Brillouin zone (BZ) taking two sites horizontally as a unit cell. At half filling, a positive $V$ drives the system from a Dirac semi-metal (DSM) to a $\mathbb{Z}_{2}$-symmetry spontaneously broken charge density wave (CDW) state, where one sublattice acquires a higher density than the other. The QCP is described by the “chiral Ising” Gross-Nevey-Yukawa (GNY) theory with global symmetry $O(N_{f})^{2}\rtimes\mathbb{Z}_{2}$ and in our case $N_{f}=2$ two-component Dirac fermions. The position of the QCP, the critical exponents of the Gross-Neveu transition, have been systematically worked out using QMC, field-theoretical expansions, and conformal bootstrap, for both $N_{f}=2$ [47, 46, 48, 49, 50, 51], and other values of $N_{f}$ [52, 46, 48, 49, 53, 50, 51]. The schematic $T-V$ phase diagram is shown in lower part of the panel (b) in Fig. 2.
@@
By simulating this model with DQMC, we can probe entanglement near the chiral GNY-Ising QCP at $V_{c}=1.27$, which we located via finite-size scaling and using the critical exponents from conformal bootstrap [46]. The detailed QMC finite-size data analysis and collapse are shown in the SM [38].
Logarithmic negativity
We present the results for $\mathcal{E}^{b}$ of the TFIM and $\mathcal{E}^{f}$ of the $t-V$ model.
Ising model— Fig. 2 (a) shows nearest neighbor $\mathcal{E}^{b}$ at finite temperature for the square lattice TFIM of linear system size $L=40$ with inverse temperature $\beta$ up to 40. The $\mathcal{E}^{b}$ maxima, marked by the purple dashed line, shift from $h>h_{c}$ at high temperature to $h\approx h_{c}=3.044$ at $T=1/40$. These numerical values align with prior studies that compute negativity as a function of $h$ rather than temperature or separation [54]. Going beyond the numerical results, we provide an argument for the absence of sudden death at small or large $h$ based on evolution of the RDM in space of states [22]. At $h=\infty$, the GS becomes the pure product state where all spins point up. The RDM of two adjacent spins is then also a pure product state, which lies at boundary of all separable (un-entangled) RDMs. The entanglement at finite $h$ is only lost at $h=\infty$. In the opposite limit $h\to 0$, the GS becomes the cat or GHZ state $(|\!\uparrow\cdots\uparrow\rangle+|\!\downarrow\cdots\downarrow\rangle)/\sqrt{2}$, leading to a rank-deficient RDM with two eigenvalues being 1/2, and the two others zero. A non-full rank separable state lies at the boundary of all separable states (see [55] and references therein), so that the RDM becomes separable only when $h$ exactly vanishes.
@@
At all $h$ values, $\mathcal{E}^{b}$ has a sudden-death temperature, beyond which $\mathcal{E}^{b}$ is identically zero corresponding to the dark blue region in the phase diagram; it is also shown in panel (c) of Fig. 3. The existence of a thermal sudden death follows from the general fact that the infinite temperature state lies well within the continent of separable states, and so entanglement must disappear at a finite temperature corresponding to the point at which the system enters the continent [22].
@@
Let us now examine the fate of entanglement as the 2 sites become progressively more separated. Fig. 3 (a) shows the $\mathcal{E}^{b}$ between two sites separated with bond-distance $r=|i-j|$. $\mathcal{E}^{b}$ drops to 0, within machine precision, even for the next-nearest neighbor separation, $r=\sqrt{2}$. Strikingly, the 2d TFIM possesses the shortest possible range of entanglement for two spins, as it only exists for nearest neighbor, even shorter than the 1d case where the sudden-death distance was found to be $r_{\textrm{sd}}=3$ [25, 26, 28, 38]. Such a sudden death separation again follows since the state at large separations lies within the separable continent [22]. Early 1d TFIM studies [26] similarly show rapid concurrence decay with temperature and distance, consistent with our results in Fig. 3(a) and (c). Additionally, Roscilde et al. [56] used quantum Monte Carlo simulations on another Ising model, confirming that concurrence remains short-ranged at a quantum phase transition.
@@
We also computed the third order moment $\mathcal{E}^{b}_{3}$ at the Ising QCP. The moments of the LN are defined [19, 20] as $\mathcal{E}^{b}_{n}=\ln\frac{\operatorname{Tr}\rho_{A}^{n}}{\operatorname{Tr}[(\rho_{A}^{T_{1}})^{n}]},$ which can be easily computed from the full RDM, or one can sample $\mathcal{E}^{b}_{n}$ by using standard QMC and exploiting the replica trick [57]. The orange line in the inset of Fig. 3 (a) shows the power-law decay of $\mathcal{E}^{b}_{3}$ with fitted power 3.9(2), which agrees with the CFT prediction $4\Delta_{\sigma}+2\approx 4.073$ [21, 45]. The power-law decay of $\mathcal{E}^{b}_{3}$ is in sharp contrast with the rapid sudden death of $\mathcal{E}^{b}$, and arises since the former is not a proper entanglement measure, and is thus polluted by non-entangling correlations.
@@
Fermion $t-V$ model.—Fig. 2 (b) shows the analogous results for $\mathcal{E}^{f}$ of two adjacent sites for the square lattice $t-V$ model of linear system size $L=6$ with $\beta$ up to 15. As denoted by the purple dashed line, the maxima starts in the ordered CDW phase at high temperature and converges to $V\approx 1$ at low temperature, which is near the QCP but in the DSM phase. It roughly follows the finite-temperature phase boundary, obtained from finite-size scaling [38] and indicated by the red line, but clearly differs from it at most temperatures. We observe that the nearest-neighbor fermionic LN is less sensitive to the QCP compared to the bosonic LN in the square lattice Ising model. A key difference is that the DSM phase has gapless Dirac fermions, and these lead to a large $\mathcal{E}^{f}$. The QCP, being driven by the fluctuations of the bosonic order parameter, thus leaves a weaker imprint on $\mathcal{E}^{f}$.